Solve for y(t) using Laplace Transformation. [20 Points] y(t) - y(t) = 0 when y(0) = 0, y(0) = 1, y(0) = 0, y(0) = 0
Answer #1

differential equation is Solution: - Giwn y) - Y11) = Taking the laplace Juans form m on on both sides b 10[154ų – $3410) –Taking inverse Laplace transform on both sides, we get :(7)=1[+(3+2+5) Y!=) = 4{(64) + 2(34;)] sint - sin ht] 1467) = 3 (s